Circular No. 5728

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET FORBES (1993f) M. P. Candy, Perth Observatory, reports the recovery of this comet

from measurements on plates taken by G. Lowe with the 0.33-m astrograph:

1993 UT  α_{2000}  δ_{2000}  m_{1}
Mar. 21.84583  21^{h}27^{m}52.^{s}66  −18^{o}45′17′′.1  14
21.88750  21 27 59.93  −18 44 37.2
27.85729  21 46 38.68  −16 58 02.3

On the first night the comet had a central condensation 5′′ across and a 1′ tail to the southwest. The indicated correction to the prediction on MPC 16379 is ΔT = +0.04 day.

SUPERNOVA 1993I IN MCG +02-32-144 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of

a possible supernova of mag 18.0 at α = 12^{h}32^{m}10^{s}.50, δ = +9^{o}16′39′′.2 (equinox 1950.0), 13′′.8 west and 8′′.6 south of the center of the S0 galaxy (B = 14.9) MCG +02-32-144. The object was found on films taken by D. Albanese and himself on Mar. 23.03 and 25.01 UT. It was also present at mag 21 on Mar. 17.02 UT but is not visible in the Palomar Sky Survey. SN 1960B also appeared in this galaxy.

GRO J0422+32 E. T. Harlaftis, Royal Greenwich Observatory; and P. A. Charles, Uni-

versity of Oxford, report: “CCD photometry of the GRO J0422+32 optical counterpart (IAUC 5588) with the 1-m Kapteyn Telescope at La Palma on Feb. 28.93 UT gave U = 15.5, B = 16.2, V = 15.3 and R = 14.8 (errors 0.2 mag) and on Mar. 19.92 UT gave V = 15.9 and R = 15.2, indicating a decay rate in V of ∼ 0.03 mag/day. A 20-min optical spectrum on Feb. 19.92 with ISIS on the 4.2-m Herschel Telescope covered the ranges 420-500 and 600-700 nm at a resolution of 0.14 nm. The main difference from the spectrum reported on IAUC 5589 is the disappearance of the 464-465-nm Bowen blend. He II (468.6 nm) is still very strong (equivalent width 2 nm) and indicates that the source is probably still x-ray active. Hα, Hβ and Hγ all show a narrow emission feature and other complex structure, with Hβ and Hγ both being superposed on a broad, shallow absorption feature (FWZI ∼ 9 nm). The 443-nm and 661.3-nm diffuse interstellar features imply E(B − V) = 0.20.”

1993 March 28 Brian G. Marsden